OpenStudy (anonymous):

solve the system of equations by the addition method 4x-7y=16 5x+37=-27

OpenStudy (anonymous):

2 Q is -12.8 and 1 Q still working on it

OpenStudy (anonymous):

i thank x=5.75 for 1Q

OpenStudy (anonymous):

y=-188/183, x= 403/183